OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -1 /2 and b = 1

OpenStudy (anonymous):

y = mx+b is the slope intercept form. how would you convert that so that it reads as a line in standard form?

OpenStudy (saifoo.khan):

y = mx + c

OpenStudy (saifoo.khan):

y = -1x/2 + 1 y = -x + 2 ------ 2 2y = -x + 2